# Branches, Assemble!  Multi-Branch Cooperation Network for Large-Scale Click-Through Rate Prediction at Taobao

### v1

@@ Abstract.
Existing click-through rate (CTR) prediction works have studied the role of feature interaction through a variety of techniques. Each interaction technique exhibits its own strength, and solely using one type could constrain the model’s capability to capture the complex feature relationships, especially for industrial large-scale data with enormous users and items. Recent research shows that effective CTR models often combine an MLP network with a dedicated feature interaction network in a two-parallel structure. However, the interplay and cooperative dynamics between different streams or branches remain under-researched. In this work, we introduce a novel Multi-Branch Cooperation Network (MBCnet) which enables multiple branch networks to collaborate with each other for better complex feature interaction modeling. Specifically, MBCnet consists of three branches: the Expert-based Feature Grouping and Crossing (EFGC) branch that promotes the model’s memorization ability of specific feature fields, the low rank Cross Net branch and Deep branch to enhance both explicit and implicit feature crossing for improved generalization. Among branches, a novel cooperation scheme is proposed based on two principles: branch co-teaching and moderate differentiation. Branch co-teaching encourages well-learned branches to support poorly-learned ones on specific training samples. Moderate differentiation advocates branches to maintain a reasonable level of difference in their feature representations. The cooperation strategy improves learning through mutual knowledge sharing via co-teaching and boosts the discovery of diverse feature interactions across branches. Extensive experiments on large-scale industrial datasets and online A/B test at Taobao app demonstrate MBCnet’s superior performance, delivering a 0.09 point increase in CTR, 1.49% growth in deals, and 1.62% rise in GMV. Core codes will be released soon.
multi-branch cooperation, CTR prediction, feature grouping, co-teaching, moderate differentiation
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06ccs: Information systems Retrieval models and ranking
@@ 1. Introduction
(a) Number of users and items
(b) Category Distribution
(c) Feature Space Coverage
Figure 1. (a) shows the data growth of users and items for Taobao’s image2product search over a two-year period. (b) illustrates the category distribution of top-10k most accurately learned samples (i.e., those with low logloss) across different branches in our MBCnet. (c) gives an example to illustrate the feature space for CTR prediction and the coverage of different approaches.
@@
Click-Through Rate (CTR) prediction which estimates the probability of a user clicking on a candidate item, is a fundamental task in online services like recommendation, retrieval, and advertising (Cheng et al., 2016; Zhou et al., 2018; Ouyang et al., 2019). The precision in predicting CTR not only substantially impacts user engagement but also has a significant effect on the revenue of industrial businesses. One key aspect of developing precise CTR prediction is to capture complex feature interactions, enabling the model to generalize across various scenarios.
@@
For many years, researchers have examined the role of feature interaction in CTR by employing a range of methodologies. Earliest algorithm is Logistic Regression (LR) (Kumar et al., 2015), which is a linear model and heavily relies on hand-crafted input features by domain experts. Factorization Machine (FM) (Rendle, 2010) combines polynomial regression models with factorization techniques to automatically capture effective feature interactions. These methods primarily focus on low-order feature interactions and have shown limitations in capturing complex patterns among features (Yang and Zhai, 2022). Due to the superior non-linear feature learning ability of Deep Neural Networks (DNN), many scholars have investigated deep learning techniques on CTR prediction (Wang, 2020). For example, Wide & Deep (Cheng et al., 2016) consists of a jointly trained wide linear model and deep neural networks, combining the benefits of memorization and generalization for recommender systems. DeepFM (Guo et al., 2017) and xDeepFM (Lian et al., 2018) combine the power of factorization machines and deep learning, to emphasize both low- and high-order feature interactions. DeepIM (Yu et al., 2020) effectively captures high-order interactions via a deep interaction machine module, serving as an efficient, exact high-order FM implementation. Instead of employing FM to model feature relationships, Wang et al. (2021, 2017a) introduced a novel cross network that explicitly performs feature crossing with each layer, which eliminates the need for manual feature engineering. Further, Mao et al. (2023) observed that even two MLP networks in parallel can achieve surprisingly good performance compared to many well-designed models. Thereby, they developed a simple yet strong dual MLP model to capture diverse feature interactions and this model has achieved state-of-the-art performance on public benchmarks.
@@
Each feature interaction technique brings its own advantages, and solely depending on one type may hinder the model’s potential to capture complex feature relationships. Particularly in industrial contexts with vast numbers of users and items, the data patterns can be exceedingly intricate. Relying on just one feature interaction technique is proved to be insufficiently effective (Ling et al., 2017). In Figure 1 (a), we show the number of users and items across by different data collection sizes in our image2product search [^1] It is an online service that provides product search via image uploads. at Taobao [^2] https://www.taobao.com/ app. The hundreds of millions of users and items indicate a broader spectrum of personalization while with higher data sparsity, bringing much more complex patterns and great challenges for modeling. As shown in Figure 1 (b) and (c), different feature interaction approaches have different modeling abilities and relying on one single approach cannot well cover the complex high-order feature space for CTR prediction. Recent researchers have highlighted that existing successful CTR models (Cheng et al., 2016; Wang et al., 2021, 2017a; Guo et al., 2017; Lian et al., 2018; Song et al., 2019) usually adopt a two-branch architecture and work in an ensemble style. The ensemble style permits the model to learn feature interactions from different perspectives. Despite the success of above models, the interplay and cooperative dynamics between different streams or branches remain under-researched. This motivates us to study the working principles of multi-branch CTR networks and develop a more effective model to capture the complex patterns.
@@
In this work, we propose a novel Multi-Branch Cooperation Network (MBCnet) which enables multiple branch networks to collaborate with each other for better complex feature interaction modeling. MBCnet consists of three different prediction branches and a cooperation scheme. One important prediction branch is a designed Expert-based Feature Grouping and Crossing (EFGC) module, which groups hundreds of feature fields and only conducts feature interaction between specific group pairs by domain-expert knowledge. This branch enhances the model’s memorization ability to capture expert-driven feature interactions for CTR prediction. The other two branches are the existing popular and powerful low-rank Cross Net (Wang et al., 2021) for explicit feature crossing and Deep Net (Wang et al., 2017a; Guo et al., 2017) for implicit feature crossing, which both improve the model’s generalization ability (Cheng et al., 2016). These three branches exhibit diverse characteristics and modeling capabilities, which can be integrated into a more powerful model. Further, to better utilize the advantages of different branches, we propose a multi-branch cooperation scheme based on two principles. The first principle (branch co-teaching) dictates that well-learned branch should assist poorly-learned branch on particular samples, while the second principle (moderate differentiation) maintains that latent features of distinct branches should preserve a moderate level of differentiation, avoiding extremes of either excessive divergence or excessive similarity. Specifically, the first principle is embodied in a co-teaching objective function, where the well-learned branch uses its predictions as pseudo labels to guide the poorly-learned one on disagreed predicted samples. The second principle is formulated as an equivalent transformation regularization loss between latent features of branches. The cooperative strategy promotes learning abilities of different branches through knowledge transfer via co-teaching and also enhances their capabilities to uncover various patterns of feature interactions.
@@
Through experiments, we have shown that MBCnet can largely improve CTR prediction performance. We have conducted an online A/B test in image2product search at Taobao app, which revealed that MBCnet could achieve an absolute 0.09 point CTR improvement, along with a relative 1.49% deal growth and a relative 1.62% GMV rise. To sum up, the contributions are summarized as follows:
@@
  • We propose a novel multi-branch cooperation network (MBCnet) which ensembles different feature interaction branches to enlarge the capacity of capturing complex feature patterns within large-scale data. Specifically, MBCnet consists of a designed EFGC branch and two popular and powerful branches. Each of them possess unique capabilities and strengths, enabling them to form a more powerful CTR prediction model.
@@
  • We introduce a novel cooperation scheme that facilitates cooperation and complements of branches to promote the overall learning ability. The cooperation scheme is constructed upon two principles: branch co-teaching and moderate differentiation. These principles promote each branch’s learning ability, while also enabling them to uncover distinct patterns even with the same supervision.
@@ 2. Related Work
Click-Through Rate (CTR) Prediction: Capturing the complex feature interactions is a key to a successful CTR model. Early works combine Logistic Regression (LR) (Kumar et al., 2015) and Factorization Machine (FM) (Rendle, 2010) to capture feature interactions. While these methods either rely much on handcrafted features or focus on lower-order feature interactions, showing limitations in learning complex feature patterns (Yang and Zhai, 2022). Leveraging the superior capacity of deep neural networks (DNN) for feature extraction, deep learning techniques have been widely examined in the context of CTR prediction (Cheng et al., 2016; Guo et al., 2017; Wang et al., 2017b). For example, Cheng et al. (2016) combined shallow linear models and deep non-linear networks to promote the model’s memorization and generalization ability for recommendation systems. Guo et al. (2017) and Lian et al. (2018) combines the power of factorization machines and deep learning for feature learning. Wang et al. (2021) designed Deep and Cross Network (DCN) which can more efficiently and explicitly learn certain bounded-degree feature interactions. They later introduced a mixture of low-rank expert network into DCN to make the model more practical in large-scale industrial settings (Wang et al., 2017a). In FinalMLP (Mao et al., 2023), researchers observed that even two parallel MLP networks can achieve satisfied performance. They further proposed feature gating and interaction aggregation layers that can be easily plugged to make an enhanced two-stream MLP model. Some other CTR works investigate sequential user behavior modeling (Zhou et al., 2018, 2019; Ouyang et al., 2019; Chen et al., 2019), cross-domain knowledge transfer (Ouyang et al., 2020; Li and Tuzhilin, 2020; Chen et al., 2022, 2020, 2024; Su et al., 2020) and multi-task learning (Ma et al., 2018; Tang et al., 2020) for improved performance. These topics are beyond the scope of this paper, so we do not elaborate them here.
@@
Existing works have studied different ways to model feature interactions and they usually contain two parallel branches for enhanced CTR performance. Our MBCnet also works in a multi-branch style, but it additionally contains a novel cooperation scheme that enables different branches to learn from each other and uncover various feature interaction patterns. Moreover, we design the EFGC branch, which can better incorporate domain-expert knowledge and promote the model’s memorization ability.
@@
Ensemble Learning: Ensemble learning is a machine learning technique that aggregates two or more learners (e.g., regression models, neural networks) to produce better prediction performance (Yang et al., 2022; Mienye and Sun, 2022). The fundamental idea is that single machine learning model suffers from various limitations (e.g., high variance, high bias, low accuracy). However, combining different models can address these limitations and achieve higher accuracy (Brown, 2010). For instance, bagging (e.g., random forest) reduces variance without increasing the bias, while boosting (e.g., XGBoost) reduces bias (Doroudi, 2020). Generally speaking, an ensemble learner is more robust and able to perform better than the individual learners. In CTR prediction, different feature interaction techniques exhibit different modeling advantages. It is necessary to combine them together to better capture the complex patterns. Researchers have investigated the ensemble idea in CTR prediction from several aspects (Ling et al., 2017). He et al. (2014) utilized GBDT for non-linear feature transformation and feed them to the LR for final prediction. While Zhang et al. (2016) used the FM embedding learned from sparse features to DNN. Researchers have found that simply replicating them cannot not yield satisfied performance. Ling et al. (2017) investigated eight ensemble variants of GBDT, LR and DNN. They found that initializing the GBDT sample target with DNN’s prediction score yielded the best performance in their business context. These methods do not work in an end-to-end style and require hard efforts on hand-craft tuning. Later, deep learning based ensemble CTR works (Cheng et al., 2016; Guo et al., 2017; Wang et al., 2017b) are becoming more popular. EnKD (Zhu et al., 2020) specifically studies knowledge distillation technique for model ensemble in order to learn light and accurate student prediction model.
Compared to EnKD (Zhu et al., 2020), our MBCnet has one novel EFGC branch for feature interaction modeling and a completely different ensemble scheme. Detailed comparisons would be discussed later.
@@
Figure 2. (a) shows the architecture of our multi-branch cooperation networks (MBCnet). It consists of three different branches: the Expert-based Feature Grouping and Crossing (EFGC) branch, the low-rank Cross Net branch and the Deep net branch. To enlarge the model’s capability, there is a novel multi-branch cooperation scheme based on two principles: branch co-teaching and moderate differentiation. (b) illustrates the branch co-teaching approach between any two of the three branches. (c) demonstrates that the latent features from any two branches undergo equivalent transformation.

### v2

# Learning Multi-Branch Cooperation for Enhanced Click-Through Rate Prediction at Taobao

@@ Abstract.
Existing click-through rate (CTR) prediction works have studied the role of feature interaction through a variety of techniques. Each interaction technique exhibits its own strength, and solely using one type usually constrains the model’s capability to capture the complex feature relationships, especially for industrial data with enormous input feature fields. Recent research shows that effective CTR models often combine an MLP network with a dedicated feature interaction network in a two-parallel structure. However, the interplay and cooperative dynamics between different streams or branches remain under-researched. In this work, we introduce a novel Multi-Branch Cooperation Network (MBCnet) which enables multiple branch networks to collaborate with each other for better complex feature interaction modeling. Specifically, MBCnet consists of three branches: the Extensible Feature Grouping and Crossing (EFGC) branch that promotes the model’s memorization ability of specific feature fields, the low rank Cross Net branch and Deep branch to enhance explicit and implicit feature crossing for improved generalization. Among these branches, a novel cooperation scheme is proposed based on two principles: Branch co-teaching and moderate differentiation. Branch co-teaching encourages well-learned branches to support poorly-learned ones on specific training samples. Moderate differentiation advocates branches to maintain a reasonable level of difference in their feature representations on the same inputs. This cooperation strategy improves learning through mutual knowledge sharing via co-teaching and boosts the discovery of diverse feature interactions across branches. Extensive experiments on large-scale industrial datasets and online A/B test at Taobao app demonstrate MBCnet’s superior performance, delivering a 0.09 point increase in CTR, 1.49% growth in deals, and 1.62% rise in GMV. Core codes are available online.
multi-branch cooperation, CTR prediction, feature grouping, branch co-teaching, moderate differentiation
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/2018/06ccs: Information systems Retrieval models and ranking
(a) Category Distribution
(b) Feature Space Coverage
Figure 1. (a) illustrates the category distribution of top-10k most accurately learned samples (i.e., those with low logloss) across different feature interaction techniques in our MBCnet. (b) gives an example to illustrate the feature space for CTR prediction and the coverage of different approaches.
@@ 1. Introduction
Click-Through Rate (CTR) prediction which estimates the probability of a user clicking on a candidate item, is a fundamental task in online services like recommendation, retrieval, and advertising (Cheng et al., 2016; Zhou et al., 2018; Ouyang et al., 2019). The precision in predicting CTR not only substantially impacts user engagement but also has a significant effect on the revenue of industrial businesses. One key aspect of developing precise CTR prediction is to capture complex feature interactions, which enables the model to generalize across various scenarios.
@@
For many years, researchers have examined the role of feature interaction in CTR by employing a range of methodologies. Earliest algorithm is Logistic Regression (LR) (Kumar et al., 2015), which is a linear model and heavily relies on hand-crafted input features by domain experts. Factorization Machine (FM) (Rendle, 2010) combines polynomial regression models with factorization techniques to automatically capture effective feature interactions. These methods primarily focus on low-order feature interactions and have shown limitations in capturing complex patterns among features (Yang and Zhai, 2022). Due to the superior non-linear feature learning ability of Deep Neural Networks (DNN), many scholars have investigated deep learning techniques on CTR prediction (Wang, 2020). For example, Wide & Deep (Cheng et al., 2016) consists of a jointly trained wide linear model and deep neural networks, combining the benefits of memorization and generalization for recommender systems. DeepFM (Guo et al., 2017) and xDeepFM (Lian et al., 2018) combine the power of factorization machines and deep learning, to emphasize both low- and high-order feature interactions. Instead of employing FM to model feature relationships, Wang et al. (2021, 2017a) introduced a novel cross network that explicitly performs feature crossing with each layer, which eliminates the need for manual feature engineering. Further, Mao et al. (2023) observed that even two MLP networks in parallel can achieve surprisingly good performance compared to many well-designed models. Thereby, they developed a simple yet strong dual MLP model to capture diverse feature interactions and this model has achieved state-of-the-art performance on public benchmarks.
@@
Each feature interaction technique brings its own advantages, and solely depending on one may hinder the model’s potential to capture complex feature relationships. Particularly in industrial contexts with vast numbers of users and items, the data patterns can be exceedingly intricate. We usually incorporate hundreds of input feature fields to learn for the data patterns, but such many feature fields also exhibit enormous useful feature interactions. The modeling ability of single feature interaction technique is limited and insufficiently effective (Ling et al., 2017). We analyze the difference of different feature interaction techniques on our industrial data. Figure 1 (a) shows the category distribution of the top-10k most accurately predicted samples (i.e., those with low logloss) across different feature interaction techniques in MBCnet. Figure 1 (b) illustrates the feature space for CTR prediction and the coverage of different branches. Through these figures, we see different feature interaction approaches have different modeling abilities. Relying on one single approach cannot well cover the complex high-order feature space for CTR prediction. Recent researchers have highlighted that existing successful CTR models (Cheng et al., 2016; Wang et al., 2021, 2017a; Guo et al., 2017; Lian et al., 2018; Song et al., 2019) usually adopt a two-branch architecture and work in an ensemble style. The ensemble style permits the model to learn feature interactions from different perspectives. Despite the success of above models, the interplay and cooperative dynamics between different streams or branches remain under-researched. Recent models ensemble different techniques for CTR prediction by prediction voting or latent feature mean pooling or concatenation, which means no explicit interplay and cooperative signals are provided. This could largely limit the whole model’s learning ability, especially when majority branches have wrong predictions. It motivates us to study the working principles of multi-branch CTR networks and develop a more effective model to capture the complex patterns.
@@
In this work, we propose a novel Multi-Branch Cooperation Network (MBCnet) which enables multiple branch networks to collaborate with each other for better complex feature interaction modeling. MBCnet consists of three different prediction branches and a cooperation scheme. One important prediction branch is a designed Extensible Feature Grouping and Crossing (EFGC) module, which groups hundreds of feature fields and only conducts feature interaction between specific group pairs by domain-driven knowledge. It is extensible to design any intended feature field interactions, and also efficient to avoid redundant feature interactions. This branch enhances the model’s memorization ability to capture domain-driven feature interactions for CTR prediction. The other two branches are the existing popular and powerful low-rank Cross Net (Wang et al., 2021) for explicit feature crossing and Deep Net (Wang et al., 2017a; Guo et al., 2017) for implicit feature crossing, which both improve the model’s generalization ability (Cheng et al., 2016). These three branches exhibit diverse characteristics and modeling capabilities, which can be integrated into a more powerful model. More importantly, we propose a multi-branch cooperation scheme based on two principles to better exert advantages of different branches. The first principle (branch co-teaching) dictates that well-learned branch should assist poorly-learned branch on particular samples, while the second principle (moderate differentiation) maintains that latent features of distinct branches should preserve a moderate level of differentiation on the same input samples, avoiding extremes of either excessive divergence or excessive similarity. Specifically, the first principle is embodied in a co-teaching objective function, where the well-learned branch uses its predictions as pseudo labels to guide the poorly-learned one on disagreed predicted samples. The second principle is formulated as an equivalent transformation regularization loss between latent features of branches. The cooperative strategy promotes learning abilities of different branches through knowledge transfer via co-teaching and also enhances their capabilities to uncover various patterns of feature interactions.
@@
Through experiments, we have shown that MBCnet can largely improve CTR prediction performance on our industrial datasets. We have conducted an online A/B test in image2product search [^1] It is Alibaba’s online service that provides product search via image uploads. at Taobao app [^2] https://www.taobao.com/, which revealed that MBCnet could achieve an absolute 0.09 point CTR improvement, along with a relative 1.49% deal growth and a relative 1.62% Gross Merchandise Value (GMV) rise. The contributions are summarized as follows:
@@
  • We propose a novel multi-branch cooperation network (MBCnet) which ensembles different feature interaction branches to enlarge the capacity of capturing complex feature patterns in industrial data. Specifically, MBCnet consists of a designed EFGC branch and two popular and powerful branches. Each of them possess unique capabilities and strengths, enabling them to form a more powerful CTR prediction model.
@@
  • We introduce a novel cooperation scheme that facilitates cooperation and complements of branches to promote the overall learning ability. The cooperation scheme is constructed upon two principles. These principles promote each branch’s learning ability, while also enabling them to uncover distinct patterns even with the same inputs.
@@ 2. Related Work
Click-Through Rate (CTR) Prediction: Capturing the complex feature interactions is a key to a successful CTR model. Early works combine Logistic Regression (LR) (Kumar et al., 2015) and Factorization Machine (FM) (Rendle, 2010) to capture feature interactions. While these methods either rely much on handcrafted features or focus on lower-order feature interactions, showing limitations in learning complex feature patterns (Yang and Zhai, 2022). With the superior capacity of deep neural networks (DNN) for feature extraction, deep learning techniques have been widely examined in the context of CTR prediction (Cheng et al., 2016; Guo et al., 2017; Wang et al., 2017b). For example, Cheng et al. (2016) combined shallow linear models and deep non-linear networks to promote the model’s memorization and generalization ability for recommendation systems. Guo et al. (2017) and Lian et al. (2018) combines the power of factorization machines and deep learning for feature learning. DeepIM (Yu et al., 2020) effectively captures high-order interactions via a deep interaction machine module, serving as an efficient, exact high-order FM implementation. Wang et al. (2021) designed Deep and Cross Network (DCN) which can more efficiently and explicitly learn certain bounded-degree feature interactions. They later introduced a mixture of low-rank expert network into DCN to make the model more practical in large-scale industrial settings (Wang et al., 2017a). In FinalMLP (Mao et al., 2023), researchers observed that even two parallel MLP networks can achieve satisfied performance. They further proposed feature gating and interaction aggregation layers that can be easily plugged to make an enhanced two-stream MLP model. Some other CTR works investigate sequential user behavior modeling (Zhou et al., 2018, 2019; Ouyang et al., 2019), cross-domain knowledge transfer (Ouyang et al., 2020; Li and Tuzhilin, 2020; Chen et al., 2022, 2020, 2024; Su et al., 2020), multi-task learning (Ma et al., 2018; Tang et al., 2020) and current popular large language models (LLMs) augmented CTR methods (Geng et al., 2024; Lin et al., 2024; Muhamed et al., 2021; Li et al., 2023; Huan et al., 2024) for improved performance. These topics are beyond the scope of this paper, so we do not elaborate them here.
@@
Existing works have studied different ways to model feature interactions and they usually contain two parallel branches for enhanced CTR performance. Our MBCnet also works in a multi-branch style, but it additionally contains a novel cooperation scheme that enables different branches to learn from each other and uncover various feature interaction patterns. Moreover, we design the EFGC branch, which can extensibly better incorporate various domain-driven knowledge and promote the model’s memorization ability.
@@
Ensemble Learning: Ensemble learning is a machine learning technique that aggregates two or more learners (e.g., regression models, neural networks) to produce better prediction performance (Yang et al., 2022; Mienye and Sun, 2022). The fundamental idea is that single machine learning model suffers from various limitations (e.g., high variance, high bias, low accuracy). However, combining different models can address these limitations and achieve higher accuracy (Brown, 2010). For instance, bagging (e.g., random forest) reduces variance without increasing the bias, while boosting (e.g., XGBoost) reduces bias (Doroudi, 2020). Generally speaking, an ensemble learner is more robust and able to perform better than the individual learners. In CTR prediction, different feature interaction techniques exhibit different modeling advantages. It is necessary to combine them together to better capture the complex patterns. Researchers have investigated the ensemble idea in CTR prediction from several aspects (Ling et al., 2017). He et al. (2014) utilized GBDT for non-linear feature transformation and feed them to the LR for final prediction. While Zhang et al. (2016) used the FM embedding learned from sparse features to DNN. Researchers have found that simply replicating them cannot not yield satisfied performance. Ling et al. (2017) investigated eight ensemble variants of GBDT, LR and DNN. They found that initializing the GBDT sample target with DNN’s prediction score yielded the best performance in their business context. These methods do not work in an end-to-end style and require hard efforts on hand-craft tuning. Later, deep learning based ensemble CTR works (Cheng et al., 2016; Guo et al., 2017; Wang et al., 2017b; Zhu et al., 2020) are becoming more popular. For instance, CETnet (Liu et al., 2024) focuses on the feature fusion of multiple models by a confidence based approach. EnKD (Zhu et al., 2020) specifically studies knowledge distillation technique for model ensemble in order to learn light and accurate student prediction model. In addition, some multi-expert models (Ma et al., 2018; Tang et al., 2020) can also be regarded as using multi-branches for CTR modeling. They typically use several parallel MLP networks for generalized feature learning.
Compared to recent ensemble works (Cheng et al., 2016; Zhu et al., 2020; Mao et al., 2023; Wang et al., 2017b) and multi-expert models (Ma et al., 2018; Tang et al., 2020), our MBCnet has one novel EFGC branch for extensible domain-driven feature interaction modeling and a completely different ensemble scheme. It combines multiple branches whose working scheme and architectures are different, while multi-expert models usually have the same expert architectures and rely on parameter initialization for learning different pattern. Furthermore, the cooperation scheme explicitly advocates the branches to learn from each other and uncover different feature patterns even under the same supervision.
@@
Figure 2. (a) shows the architecture of our multi-branch cooperation networks (MBCnet). It consists of three different branches: the EFGC branch, the low-rank Cross Net branch and the Deep net branch. To enlarge the model’s capability, there is a novel multi-branch cooperation scheme based on two principles: branch co-teaching and moderate differentiation. (b) illustrates the branch co-teaching approach between any two of the three branches. (c) demonstrates that latent features from any two branches undergo equivalent transformation.